OpenStudy (anonymous):

4(4x + 1) = 3x - 4(6 - x) + 8 i keep getting -4 :(

OpenStudy (anonymous):

x=-20/9

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

thanks idk what i was doing wrong

OpenStudy (anonymous):

Do you still have your working?

OpenStudy (anonymous):

work?

OpenStudy (anonymous):

I just wanted to know whether or not you understood how he got -20/9.

OpenStudy (anonymous):

Now I do, but this one I am having trouble with. 3(95 - 6x) = x - 152

OpenStudy (anonymous):

What answer did you get for that one? ^

OpenStudy (anonymous):

x=23

OpenStudy (anonymous):

That's correct!

OpenStudy (anonymous):

oh sweet thanks for being awesome & helping me out

OpenStudy (anonymous):

No problem!